OpenStudy (anonymous):

how can you rewrite -3m to the power of -2 using positive exponents

OpenStudy (anonymous):

\[-3m ^{-2}\]

OpenStudy (raden):

use the property : a^(-k) = 1/a^k

OpenStudy (anonymous):

thanks

OpenStudy (raden):

welcome

OpenStudy (anonymous):

-4z^-4 simplify

OpenStudy (raden):

is that -4z^-4 or (-4z)^-4 same question for number 1 :)

OpenStudy (anonymous):

\[-4z ^{-4}\]

OpenStudy (anonymous):

i need the answer

OpenStudy (raden):

-3m^-2 = -3 * 1/m^2 = -3/m^2 -4z^-4 = -4 * 1/z^4 = -4/z^4

OpenStudy (anonymous):

\[\frac{ 1 }{ ?-3x ^{-2} }\]

OpenStudy (raden):

-1/3 * x^2

OpenStudy (anonymous):

\[2x ^{-5}\]

OpenStudy (raden):

= 2/x^5

OpenStudy (anonymous):

\[-2m ^{4}\]